OpenStudy (anonymous):

which of the following are quadratic equations A.x^3-5x^2+11x=0 B.x^2-x-2=0 C.x-5=0 D.2x^2+7=0 E.2x^2sqroot6x+8.9=0

OpenStudy (mathstudent55):

A quadratic equation must have at least one term of second degree, and no terms with higher degree. A second degree term is either a single variable that is squared, such as y^2 or x^2, or a product of two variables to the first power, like xy.

OpenStudy (mathstudent55):

It can't have square roots of variables, or variables in the denominator.

OpenStudy (anonymous):

can u help me get the answers ?

OpenStudy (mathstudent55):

Look at each choice. A. There is a term x^3. That term is of degree 3, which is a higher degree than 2. That is not quadratic.

OpenStudy (mathstudent55):

Now let's look at B. There is a 2nd degree term, x^2. There are no higher degree terms. There is no variable in the denominator or in a square root. It is quadratic.

OpenStudy (mathstudent55):

Now you tell me what you think of C.

OpenStudy (anonymous):

its a quadratic

OpenStudy (anonymous):

sorry its Not a quadratic D is a quadratic am i correct ?

OpenStudy (mathstudent55):

Do you see a second degree term in choice C? In C, is there a product of variables or a variable squared?

OpenStudy (anonymous):

the answers are B and D is that correct ?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

can u help me some more

OpenStudy (anonymous):

assume that the area of the blue region is 100 units squared. which equation represents it A.x^2+15x-50=0 B.x^2+15x+50=0 C.x^2-15x+50=0 D.x^2-15x-50=0

OpenStudy (mathstudent55):

Is there a figure?

OpenStudy (anonymous):